Diffusion Morning Report 09/06/11

Ted Berg tberg at stanford.edu
Tue Sep 6 07:28:57 PDT 2011


Good Morning All,
     Looking OK this morning. New issues: Tylan Nitride not pumping. 
Tylan 2 and 4 have MFS issues.
     Left over issues: Epi LLA issues, SiGe pump should have been back 
last week. TEOS needs test after ampule change. Thermco 2 possible Temp 
issues. No other issues at this time.

					
	Date: 	9/6/11 	THICKNESS-Change at 		Priority
Tube Name/Number 	Status 				
Tylan 1 	*up* 	** 	** 	** 	**
Tylan 2 	*yellow* 	*MFS alarms-Temp readout* 	** 	** 	*2*
Tylan 3 	*up* 	** 	** 	** 	**
Tylan 4 	*yellow* 	*MFS alarms* 	** 	** 	*3*
Tylan 5-BBr3 	*up* 	** 	** 	** 	**
Tylan 6-POCl3 	*up* 	** 	** 	** 	**
Tylan 7-TEOS2 	*yellow* 	*needs test* 	*5.9/30* 	** 	**
Tylan 8-SiGe 	*down* 	*Mech Pump problems-out for repair* 	*1.94/30* 
** 	**
Tylan 9-Poly 	*up* 	** 	*28.91/30* 	** 	**
Tylan 10-Nitride 	*down* 	*Wont Pump* 	*8.2/30* 	** 	*1*
Tylan 11-LTO 	*up* 	** 	*12.76/30* 	** 	**
Tylan 12-FGA 	*up* 	** 	** 	** 	**
Thermco 1 	*up* 	** 	** 	** 	**
Thermco 2 	*yellow* 	*Possible temp issues* 	** 	** 	**
Thermco 3 	*future* 	** 	** 	** 	**
Thermco4 	*future* 	** 	** 	** 	**
Thermco Poly1 	*up* 	*
* 	** 	** 	**
	*up* 	** 	** 	** 	**
Thermco Nitride1 	*up* 	** 	** 	** 	**
Thermco bpsg 	*up* 	** 	** 	** 	**
AMAT Epi 	*yellow* 	*lla issues* 	** 	** 	**
TEL Spa 	*up* 	** 	** 	** 	**
Burn Box 	*up* 	** 	** 	** 	**
Ebeam 	*down* 	*cryo compresssor dead* 	** 	** 	**
ZYGO 	*up* 	** 	** 	** 	**
FGA2 	*up* 	** 	** 	** 	**
SiH4 for furnaces 	** 	*B on line pressure 817 psi* 	*changeover working 
well* 	** 	**


-------------- next part --------------
An HTML attachment was scrubbed...
URL: <http://snf.stanford.edu/pipermail/diffusion/attachments/20110906/bd7ac5e6/attachment.html>


More information about the diffusion mailing list